A triangular window frame has a base of 4 ft and a height of 6 ft. How much glass will be needed to fill the window frame?
Answer by stanbon(50455)

A triangular window frame has a base of 4 ft and a height of 6 ft. How much glass will be needed to fill the window frame?
----
Area = (1/2)base*height
Area = (1/2)4*6 = 12 sq. ft.
===============
Cheers,
Stan H.